1.8²-×=2

2. (1/2)×←1/8

3. 5×=25×-²

4. 3×+²→8×-¹

5. 4³×=8×-¹


1
Expert's answer
2021-11-02T10:52:09-0400

1.

"8^{2-x}=2"

"2^{3(2-x)}=2^1"

"6-3x=1"

"x=\dfrac{5}{3}"

2.

"(\dfrac{1}{2})^{x}=\dfrac{1}{8}"

"(\dfrac{1}{2})^{x}=(\dfrac{1}{2})^3"

"x=3"

3.

"5^{x}=25^{x-2}"

"5^x=5^{2(x-2)}"

"x=2x-4"

"x=4"

4.

"3^{x+2}=8^{x-1}"

"\ln3^{x+2}=\ln8^{x-1}"

"(x+2)\ln3=(x-1)\ln8"

"x(\ln8-\ln3)=2\ln3+\ln8"

"x=\dfrac{2\ln3+\ln8}{\ln8-\ln3}"

5.

"4^{3x}=8^{x-1}"

"2^{2(3x)}=2^{3(x-1)}"

"6x=3x-3"

"x=-1"